Please help me understand this word problem. How could I solve this.
Find the dimensions of a right triangle if the shortest side is 3 less than the middle one and 6 less than the hypotenuse.

Answer by Nate(3500)

The middle sized side: x
Smallest size: x - 3
Hypotenuse: x - 3 + 6 or x + 3
(x + 3)^2 = x^2 + (x - 3)^2
x^2 + 6x + 9 = x^2 + x^2 - 6x + 9
0 = x^2 - 12x
0 = x(x - 12)
x = 0 ... not possible
x = 12 ... possible
Middle sized side: 12
Smallest size: 9
Hypotenuse: 15